# Fermat near misses and the integral Hilbert property

### v1

@@ 4.1. Double conic bundles
    $\mathcal{X}\setminus\mathcal{D}$ admits an ${\mathcal{O}}_{k,S}$-integral point which lies on a smooth fibre $C$ of $\pi_{1}$ such that $C^{\circ}({\mathcal{O}}_{k,S})$ is infinite and $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$ is unramified along $C\cap D$.
@@
Let $X$ be a smooth cubic surface over a field $k$ which contains two distinct lines $L_{1},L_{2}\subset X$. Consider the associated conic bundles $\pi_{1}$ and $\pi_{2}$ with divisor classes $C_{1}$ and $C_{2}$. Let $E$ be the divisor given by the common irreducible components of both fibrations and $D$ a smooth hyperplane section.
@@
Assume that (1) holds. It is well-known that two skew lines meet exactly 5 others; this can be seen from the fact that there are 10 lines which meet $L_{1}$, corresponding to the $5$ singular fibres of the conic bundle, and $L_{2}$ gives a section of the conic bundle, thus meets exactly 5 of these lines. From this description it is clear that they are pairwise skew. The complement is not simply connected by Lemma 3.2
Now assume that we are in case (2). The divisor $E$ is given by the lines which lie in a common plane with both $L_{1}$ and $L_{2}$. As $L_{1}$ and $L_{2}$ are coplanar there is a unique such line. By Lemma 3.1 it suffices to note that the complement of a smooth hyperplane section for a quartic del Pezzo surface is simply connected [Har17, Lem. 3.3.6]. ∎
@@
$\pi_{1}:(x_{0},x_{1},x_{2},x_{3})\mapstochar\rightarrow\begin{cases}(x_{0},x_{2})&\text{ if }(x_{0},x_{2})\neq(0,0)\\ (-x_{2}^{2}-3x_{3}^{2},x_{0}^{2}+3x_{1}^{2})&\text{ if }(x_{0}^{2}+3x_{1}^{2},x_{2}^{2}+3x_{3}^{2})\neq(0,0).\end{cases}$
@@
We next find an integer point on the surface which lies on a conic that meets the boundary divisor in a real quadratic point. Using the isomorphism ${\mathbb{P}}^{2}\rightarrow P_{s,t}$, $(y_{0},y_{1},y_{2})\mapstochar\rightarrow(sy_{0},y_{1},ty_{0},y_{2})$, we can write
@@
The intersection of the conic and the boundary divisor $D:y_{2}-ty_{0}=0$ is $(s^{3}+4t^{3})y_{0}^{2}+3sy_{1}^{2}=0$, For this to be real quadratic we require that $s(s^{3}+4t^{3})>0$ and is not a square. The solution (x,y,w,z)=(9,8,1,6)(x,y,w,z)=(-9,8,1,6), which gives $(x_{0},x_{1},x_{2},x_{3})=(1,17,-7,5)$ satisfies this real condition, as required: it lies on the conic $C=\pi_{1}^{-1}(1,-7)$. Furthermore, one can verify that the conic bundle induced by the second line and passing through $C_{1,-7}$ intersect $D$ in reduced points, thus it is unramified on $C_{1,-7}\cap D$.
@@ 5.1. Geometry of Fermat del Pezzo surfaces
Now assume that ${\mathbb{Q}}(\zeta,\sqrt[4]{n}):{\mathbb{Q}}]<16$. This happens if and only if $n$ becomes a square in ${\mathbb{Q}}(\zeta)$. Since $\ker({\mathbb{Q}}^{\times}/{\mathbb{Q}}^{\times 2}\rightarrow{\mathbb{Q}}(\zeta)^{\times}/{\mathbb{Q}}(\zeta)^{\times 2})=\{\pm 1,\pm 2\}$, this happens if and only if one of $\pm n,\pm 2n$ is a square.
First consider the case where n-n is a square. We subdivide into the cases depending on whether n-n becomes a square or a fourth power in ${\mathbb{Q}}(\zeta)$. Since $\ker({\mathbb{Q}}^{\times}/{\mathbb{Q}}^{\times 4}\rightarrow{\mathbb{Q}}(\zeta)^{\times}/{\mathbb{Q}}(\zeta)^{\times 4})=\{\pm 1,\pm 4\}$, we see that n-n becomes a fourth power in ${\mathbb{Q}}(\zeta)$ if and only if n-n or 4n-4n is a fourth power. However if n-n is fourth power, or n-n is a square but not a fourth power, there is no conic bundle structure, as follows from the fact that $\operatorname{rank}\operatorname{Pic}X_{-1}=\operatorname{rank}\operatorname{Pic}X_{-9}=1$.
@@
We now write down explicitly the conic bundle structures in Proposition 5.1. Let $X$ be a del Pezzo surface of degree $2$ with a conic bundle structure $X\rightarrow{\mathbb{P}}^{1}$. Combining with the anticanonical map $X\rightarrow{\mathbb{P}}^{2},(x,y,z,w)\mapstochar\rightarrow(x,y,z)$ yields an embedding $X\rightarrow{\mathbb{P}}^{1}\times{\mathbb{P}}^{2}$ as a surface of bidegree $(2,2)$. This embedding gives a clearer way to visualise the conic bundle structure. In the statement of the next proposition, the formula for the conic bundle gives a priori only a rational map; that it extends to a morphism follows from Lemma 2.11. We include the factorisation of the determinant of the corresponding quadratic form; this gives the locus of the singular fibres and is easy to calculate from the conic bundle equation.
@@
The surfaces in $X_{n}$ from Proposition 5.1 admit the conic bundle morphism $\pi:X_{n}\rightarrow{\mathbb{P}}^{1}$ with induced image in ${\mathbb{P}}^{1}\times{\mathbb{P}}^{2}$, given by Table 5.1.
Case Conic bundle, equation in ${\mathbb{P}}^{1}\times{\mathbb{P}}^{2}$ and determinant
$n=-4m^{4}$ $\pi:X_{n}\rightarrow{\mathbb{P}}^{1},\quad(x,y,z,w)\mapstochar\rightarrow(x^{2}+2mxz+2m^{2}z^{2},w-y^{2})$
@@
$n=m^{2}$ $\pi:X_{n}\rightarrow{\mathbb{P}}^{1},\quad(x,y,z,w)\mapstochar\rightarrow(x^{2}-mz^{2},w-y^{2})$
@@
$n=2m^{2}$ $\pi:X_{n}\rightarrow{\mathbb{P}}^{1},\quad(x,y,z,w)\mapstochar\rightarrow(xy-mz^{2},x^{2}-y^{2}-w)$
@@
$n=-2m^{2}$ $\pi:X_{n}\rightarrow{\mathbb{P}}^{1},\quad(x,y,z,w)\mapstochar\rightarrow(xy-mz^{2},x^{2}+y^{2}-w)$
@@
We achieve the descriptions by turning the proof of Lemma 2.11 into an effective method. Note also that $X_{n}({\mathbb{Q}})\neq\emptyset$ since they contain the rational point $(0,1,0,1)$, as necessary for Lemma 2.11. The first step will be to write down Galois invariant collections of lines which meet in pairs and are otherwise disjoint.
@@
which form the singular fibre $F=L+L^{\prime}$ of a conic bundle defined over $\mathbb{Q}$}. We calculate $\mathrm{H}^{0}(X_{n},F)$ by recalling that an element of this cohomology group can be represented by a rational function $f$ such that $\mathrm{div}f+F\geqslant 0$. This gives $\mathrm{H}^{0}(X_{n},F)=\langle 1,(w-y^{2})/(x^{2}+2mxz+2m^{2}z^{2})\rangle$, hence the conic bundle morphism is as stated in the proposition. One easily calculates that it has the given image (we found the equation using Magma).
@@
$\tau:\zeta\mapstochar\rightarrow\zeta^{3},\quad\sigma:\zeta\mapstochar\rightarrow\zeta^{5}.$ (5.2)
@@
Therefore the map $X_{n}\rightarrow{\mathbb{P}}^{2}$ induced by $|F+F^{\sigma}|$ is
@@
${\mathbb{P}}^{2}\dashrightarrow{\mathbb{P}}^{1},\quad(x_{0},x_{1},x_{2})\mapstochar\rightarrow(2x_{0}-x_{1},x_{2}).$ (5.4)
@@
$\begin{matrix}X_{n}&\dashrightarrow&{\mathbb{P}}^{1}\\ (x,y,z,w)&\mapstochar\rightarrow&(w^{2}+2m^{2}z^{4}+wy^{2}-wx^{2}-2mxyz^{2},2(wxy-m(x^{2}-y^{2})z^{2})).\end{matrix}$
@@
$\begin{matrix}X_{n}&\dashrightarrow&{\mathbb{P}}^{1}\\ (x,y,z,w)&\mapstochar\rightarrow&(w^{2}-2m^{2}z^{4}+2mxyz^{2}-wx^{2}-wy^{2},2(m(x^{2}+y^{2})z^{2}-wxy)).\end{matrix}$
@@
By Theorem 1.3, it suffices to show that our conditions correspond to an integral point which lies on a smooth conic that meets the boundary divisor in a real quadratic point and such that the dual conic bundle is unramified. We use Proposition 5.2, which gives explicit descriptions for the conic bundle structures for the surfaces in Theorem 1.1, which we denote by $\pi_{1}$ with dual conic bundle $\pi_{2}$ obtained via the Geiser involution. For n=1,4n=1,-4 one checks that the points given in Table 1.1 have this property. So we assume for the remainder of the proof that we are not in any of these cases, up to a fourth power.
@@
    If a smooth fibre $C$ for $\pi_{1}$ meets $D$ in a reduced point, then $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$ is unramified along $C\cap D$.
@@
(3) Our boundary divisor $D$ has the special property that it is preserved under the action of the Geiser involution. It follows that a smooth fibre $C_{1}$ for $\pi_{1}$ meets $D$ in a non-reduced point if and only if $\iota(C_{1})$ meets $D$ in a non-reduced point. The claim is now immediate. ∎
Lemma 5.4 shows that, away from the fibres over $st=0$ in the case $n=\pm 2m^{2}$, any integral point lies on a smooth conic which meets the boundary divisor in a closed point of degree two which is an unramified point with respect to $\pi_{2}$. It remains to verify that the real conditions in Theorem 1.1 are now exactly asking that the residue field of this point is real quadratic.
For the case $n=m^{2}$ this is clear from the equation. For $n=2m^{2}$, the discriminant of the field where a fibre meets the boundary divisor is $4s^{4}+12s^{2}t^{2}+t^{4}$ which is positive definite, hence the field is always real quadratic. The corresponding discriminant for the case $n=-2m^{2}$ is $4s^{4}-12s^{2}t^{2}+t^{4}$, so we require this to be positive, as in the statement of Theorem 1.1. ∎

### v2

@@ 4.1. Double conic bundles
    $\mathcal{X}\setminus\mathcal{D}$ admits an ${\mathcal{O}}_{k,S}$-integral point which lies on a smooth fibre $C$ of $\pi_{1}$ such that $C^{\circ}({\mathcal{O}}_{k,S})$ is infinite and $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$ is unramified along $C\cap D$.
@@
Let $X$ be a smooth cubic surface over a field $k$ which contains two distinct lines $L_{1},L_{2}\subset X$. Consider the associated conic bundles $\pi_{1}$ and $\pi_{2}$ with divisor classes $C_{1}$ and $C_{2}$. Let $E$ be the divisor given by the common irreducible components of both fibrations and let $D$ be a smooth hyperplane section.
@@
(1) It is well-known that two skew lines meet exactly 5 others; this can be seen from the fact that there are 10 lines which meet $L_{1}$, corresponding to the $5$ singular fibres of the conic bundle, and $L_{2}$ gives a section of the conic bundle, thus meets exactly 5 of these lines. From this description it is clear that they are pairwise skew. The complement is not simply connected by Lemma 3.2
(2) The divisor $E$ is given by the lines which lie in a common plane with both $L_{1}$ and $L_{2}$. As $L_{1}$ and $L_{2}$ are coplanar there is a unique such line. By Lemma 3.1 it suffices to note that the complement of a smooth hyperplane section for a quartic del Pezzo surface is simply connected [Har17, Lem. 3.3.6]. ∎
@@
$\pi_{1}:(x_{0},x_{1},x_{2},x_{3})\mapsto\begin{cases}(x_{0},x_{2})&\text{ if }(x_{0},x_{2})\neq(0,0)\\ (-x_{2}^{2}-3x_{3}^{2},x_{0}^{2}+3x_{1}^{2})&\text{ if }(x_{0}^{2}+3x_{1}^{2},x_{2}^{2}+3x_{3}^{2})\neq(0,0).\end{cases}$
@@
We next find an integer point on the surface which lies on a conic that meets the boundary divisor in a real quadratic point. Using the isomorphism ${\mathbb{P}}^{2}\rightarrow P_{s,t}$, $(y_{0},y_{1},y_{2})\mapsto(sy_{0},y_{1},ty_{0},y_{2})$, we can write
@@
This equation can be interpreted as lying in some ${\mathbb{P}}^{2}$-bundle over ${\mathbb{P}}^{1}$; see [Frei2018RationalSurfaces, §2] for this perspective. The intersection of the conic and the boundary divisor $D:y_{2}-ty_{0}=0$ is $(s^{3}+4t^{3})y_{0}^{2}+3sy_{1}^{2}=0$. For this to be real quadratic we require that $s(s^{3}+4t^{3})>0$ and that this is a non-square. The solution (x,y,w,z)=(9,8,1,6)(x,y,w,z)=(-9,8,1,6), which gives $(x_{0},x_{1},x_{2},x_{3})=(1,17,-7,5)$ satisfies this real condition, as required: it lies on the conic $C=\pi_{1}^{-1}(1,-7)$. As for the unramified condition in Theorem 1.4, the ramification points of $D$ with respect to $\pi_{2}$ exactly correspond to the planes through $L_{2}$ which meet $D$ in a non-reduced subscheme. However the planes passing through the points $C_{1,-7}\cap D$ and $L_{2}$ are given by equations
$(1-\sqrt{457})(x+z)+(13-\sqrt{457})(y+w)=0,\quad(1+\sqrt{457})(x+z)+(13+\sqrt{457})(y+w)=0.$
Then one checks that these planes meet $D$ in a reduced subscheme, as required. Thus Theorem 1.4 implies that the integral points are non-thin.
@@ 5.1. Geometry of Fermat del Pezzo surfaces
Now assume that ${\mathbb{Q}}(\zeta,\sqrt[4]{n}):{\mathbb{Q}}]<16$. This happens if and only if $n$ becomes a square in ${\mathbb{Q}}(\zeta)$. Since $\ker({\mathbb{Q}}^{\times}/{\mathbb{Q}}^{\times 2}\to{\mathbb{Q}}(\zeta)^{\times}/{\mathbb{Q}}(\zeta)^{\times 2})=\{\pm 1,\pm 2\}$, this happens if and only if one of $\pm n,\pm 2n$ is a square.
First consider the case where n-n is a square. We subdivide into the cases depending on whether n-n becomes a square or a fourth power in ${\mathbb{Q}}(\zeta)$. Since $\ker({\mathbb{Q}}^{\times}/{\mathbb{Q}}^{\times 4}\to{\mathbb{Q}}(\zeta)^{\times}/{\mathbb{Q}}(\zeta)^{\times 4})=\{\pm 1,\pm 4\}$, we see that n-n becomes a fourth power in ${\mathbb{Q}}(\zeta)$ if and only if n-n or 4n-4n is a fourth power. However if n-n is fourth power, or n-n is a square but not a fourth power, there is no conic bundle structure, as follows from the fact that $\operatorname{rank}\operatorname{Pic}X_{-1}=\operatorname{rank}\operatorname{Pic}X_{-9}=1$.
@@
We now write down explicitly the conic bundle structures in Proposition 5.1. Let $X$ be a del Pezzo surface of degree $2$ with a conic bundle structure $X\to{\mathbb{P}}^{1}$. Combining with the anticanonical map $X\to{\mathbb{P}}^{2},(x,y,z,w)\mapsto(x,y,z)$ yields an embedding $X\to{\mathbb{P}}^{1}\times{\mathbb{P}}^{2}$ as a surface of bidegree $(2,2)$ (see [Frei2018RationalSurfaces, §5.5.2] for more general descriptions of conic bundles on del Pezzo surfaces). This embedding gives a clearer way to visualise the conic bundle structure. In the statement of the next proposition, the formula for the conic bundle gives a priori only a rational map; that it extends to a morphism follows from Lemma 2.11. We include the factorisation of the determinant of the corresponding quadratic form; this gives the locus of the singular fibres and is easy to calculate from the conic bundle equation.
@@
The surfaces in $X_{n}$ from Proposition 5.1 admit the conic bundle morphism $\pi:X_{n}\to{\mathbb{P}}^{1}$ with induced image in ${\mathbb{P}}^{1}\times{\mathbb{P}}^{2}$, given by Table 5.1.
Case Conic bundle, equation in ${\mathbb{P}}^{1}\times{\mathbb{P}}^{2}$, and determinant
$n=-4m^{4}$ $\pi:X_{n}\to{\mathbb{P}}^{1},\quad(x,y,z,w)\mapsto(x^{2}+2mxz+2m^{2}z^{2},w-y^{2})$
@@
$n=m^{2}$ $\pi:X_{n}\to{\mathbb{P}}^{1},\quad(x,y,z,w)\mapsto(x^{2}-mz^{2},w-y^{2})$
@@
$n=2m^{2}$ $\pi:X_{n}\to{\mathbb{P}}^{1},\quad(x,y,z,w)\mapsto(xy-mz^{2},x^{2}-y^{2}-w)$
@@
$n=-2m^{2}$ $\pi:X_{n}\to{\mathbb{P}}^{1},\quad(x,y,z,w)\mapsto(xy-mz^{2},x^{2}+y^{2}-w)$
@@
We achieve the descriptions by turning the proof of Lemma 2.11 into an effective method. Note that $X_{n}({\mathbb{Q}})\neq\emptyset$ since they contain the rational point $(0,1,0,1)$, as necessary for Lemma 2.11. The first step will be to write down Galois invariant collections of lines which meet in pairs and are otherwise disjoint.
@@
which form the singular fibre $F=L+L^{\prime}$ of a conic bundle defined over $\mathbb{Q}$}. We calculate $\mathrm{H}^{0}(X_{n},F)$ by recalling that an element of this cohomology group can be represented by a rational function $f$ such that $\mathrm{div}f+F\geqslant 0$. This gives $\mathrm{H}^{0}(X_{n},F)=\langle 1,(w-y^{2})/(x^{2}+2mxz+2m^{2}z^{2})\rangle$, hence the conic bundle morphism is as stated in the proposition. One easily calculates that it has the given image (we found the equation using Magma [Magma]).
@@
$\tau:\zeta\mapsto\zeta^{3},\quad\sigma:\zeta\mapsto\zeta^{5}.$ (5.2)
@@
Therefore the map $X_{n}\to{\mathbb{P}}^{2}$ induced by $|F+F^{\sigma}|$ is
@@
${\mathbb{P}}^{2}\dashrightarrow{\mathbb{P}}^{1},\quad(x_{0},x_{1},x_{2})\mapsto(2x_{0}-x_{1},x_{2}).$ (5.4)
@@
$\begin{matrix}X_{n}&\dashrightarrow&{\mathbb{P}}^{1}\\ (x,y,z,w)&\mapsto&(w^{2}+2m^{2}z^{4}+wy^{2}-wx^{2}-2mxyz^{2},2(wxy-m(x^{2}-y^{2})z^{2})).\end{matrix}$
@@
$\begin{matrix}X_{n}&\dashrightarrow&{\mathbb{P}}^{1}\\ (x,y,z,w)&\mapsto&(w^{2}-2m^{2}z^{4}+2mxyz^{2}-wx^{2}-wy^{2},2(m(x^{2}+y^{2})z^{2}-wxy)).\end{matrix}$
@@
By Theorem 1.3, it suffices to show that our conditions correspond to an integral point which lies on a smooth conic that meets the boundary divisor in a real quadratic point and such that the dual conic bundle is unramified. We use Proposition 5.2, which gives explicit descriptions for the conic bundle structures for the surfaces in Theorem 1.1, which we denote by $\pi_{1}$ with dual conic bundle $\pi_{2}$ obtained via the Geiser involution. For n=1,4n=1,-4 one checks that the points given in Table 1.1 have this property. So we assume for the remainder of the proof that we are not in any of these cases, up to a fourth power. With respect to the conic bundle model, the boundary divisor is given by $D:x_{2}=0$. From Table 5.1, one easily deduces that the discriminant of $\pi_{1}|_{D}:D\to{\mathbb{P}}^{1}$ is given as:
$n$ $m^{2}$ $2m^{2}$ $2m^{2}$
8st(st)(s+t)8st(s-t)(s+t) $4s^{4}+12s^{2}t^{2}+t^{4}$ $(2s^{2}-4st+t^{2})(2s^{2}+4st+t^{2})$
Table 5.2. Discriminant of $\pi_{1}|_{D}:D\to{\mathbb{P}}^{1}$
@@
    For every smooth fibre $C$ for $\pi_{1}$, the map $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$ is unramified along $C\cap D$.
@@
(3) Our boundary divisor $D$ has the special property that it is preserved under the action of the Geiser involution $\iota:w\mapsto-w$. One checks that $\iota$ acts on $D$ via:
$n$ $m^{2}$ $2m^{2}$ $2m^{2}$
$\mapsto (t,s)(t,-s)$ (t,2s)(t,-2s) $(t,2s)$
Let $R_{i}\subset D$ denote the ramification locus of $\pi_{i}|_{D}:D\to{\mathbb{P}}^{1}$, viewed as a closed subscheme of degree $4$; we have $\iota(R_{1})=R_{2}$. In each case one checks that the action preserves the discriminant of $\pi_{1}|_{D}$ as in Table 5.2. It follows that $\iota$ preserves $R_{1}$, thus $R_{1}=R_{2}$. However it also follows from Tables 5.1 and 5.2 that $D$ ramifies only above the singular fibres of $\pi_{1}$, thus the same holds for $\pi_{2}$ since $\iota$ preserves intersection multiplicities and maps singular fibres to singular fibres.
Now assume for a contradiction that $C$ is a smooth fibre for $\pi_{1}$ such that $C\cap R_{2}\neq\emptyset$. It follows that $\iota(C)\cap R_{1}\neq\emptyset$. However from the above we know that it is only the singular fibres of $\pi_{2}$ which pass through $R_{1}=R_{2}$. It follows that $\iota(C)$ is singular, thus $C$ is singular, which is a contradiction. ∎
Lemma 5.4 shows that, away from the fibres over $st=0$ in the case $n=\pm 2m^{2}$, any integral point lies on a smooth conic which meets the boundary divisor in a closed point of degree two which is an unramified point with respect to $\pi_{2}$. It remains to verify that the real conditions in Theorem 1.1 are now exactly asking that the residue field of this point is real quadratic. However this is now clear from Table 5.2, noting that for $n=2m^{2}$ the discriminant $4s^{4}+12s^{2}t^{2}+t^{4}$ is positive definite, and for $n=-2m^{2}$ the discriminant expands to $4s^{4}-12s^{2}t^{2}+t^{4}$. ∎
Remark $5.$
By Lemma 5.4, the trivial solutions presented in Remark 1.2 exactly correspond to integral points which lie on a conic that meets the divisor $D$ in 22 rational points. Such conics have only finitely many integral points (Lemma 2.5), so are useless for our method.